Crew Photos – February 17th

02172017 Leaving the hab


12N 518500 4250000 02172017 Installing the seismometer
12N 518000 4250100 02172017 First steps on Mars
12N 518000 4250100 02172017 Landsacpe
12N 518000 4250100 02172017 Mouadh
12N 518000 4250100 02172017 Arthur
12N 518000 4250100 02172017 Crew exploring
12N 518700 4250700 02172017 Victoria
02172017 We now master bread

GreenHab Report – February 17th

Crew 175 GreenHab report 17Feb2017

Green Hab Report  – Sol 5
Report written by: Victoria DA-POIAN (Crew Biologist)

Date : 02/17/2017

Functionality: The heater in the green hab is working well. Today was a very cloudy and windy day ! As I was in EVA this morning, Louis Maller has checked the temperature in the GreenHab this morning. It was around 24 Celsius degrees around 9:40 AM while the temperature in the tunnel was around 10.4 Celsius degrees. He switched on the cooler (on step 1) and the GreenHab temperature was 20 Celsius degrees at 4:00PM while it was 14 Celsius degrees in the tunnel. I switched the cooler off. Louis Maller watered the seedlings this morning, and I did this afternoon too.

Status: The existing seedlings in the green hab are continuing to grow well. There are spinach, lettuce, radish, and beans growing very well in the small pots.

The lettuces are growing well too.

The Vegidair has been installed yesterday and is functionning well.

Planned activity: I have checked my new plants (planted yesterday) , the ones I want to compare with the ones I have in the Vegidair.



Journalist Report – February 17th

Crew 175 Journalist report Sol 5

Journalist report, 02/17/17 – Sol 5: Seismometer, storm and fresh bread.

Today was my first day following an EVA from the hab. Moreover, I was its habcom. After breakfast, I was in charge of checking everyone’s radio, backpack, timing the pressurisation, and giving tasks to the crew members outside during the check-ups at the beginning of the EVA. Then I was following their actions by radio when they were in range.

Today’s EVA included Arthur, as leader, Mouâdh, Victoria, and Simon, who had taken the camera, so that today’s pictures of the outside would not be mine today. Their first objective was to install the seismometer where we planned to two days ago. It went well and was finished quickly. The balloon remained in the hab, because of the wind blowing too hard. That is why, after half of the EVA, the team started to explore the surroundings, went up to hab ridge road, then back, training on how to use the map in situation.

At the same time, in the hab, I was busy working on interviews for French medias, Xavier was working on the energy supply of the station, and Louis was busy with the EMUI. The afternoon started quietly, with almost everybody taking a nap, but after a couple of hours, wind started to blow very strongly. The whole station became very noisy, parts moving everywhere, we had the strange feeling that we could take off at any minute… Our brand-new wall was moving dangerously, but went through it. Xavier and Louis had to go out on an “emergency EVA” to check if nothing was damaged, and to pick up stuff that flew away.

For dinner, Simon had baked fresh bread for us. He now masters the art of using the bread machine well. It might seem simple but it is the only way for us to get fresh food. And for French people, being able to eat fresh bread even if it isn’t baguette is priceless…


Louis MANGIN, crew journalist MDRS 175

Science Report – February 17th

Crew 175 Science Report 17Feb2017

Science Report SOL 5

Experiment: Seismometer
Person filling in the report: Mouadh Bouayad

We could deploy the seismometer this morning, during the EVA. It was quite hard to do, with the gloves, especially to put the seismometer correctly. It was nevertheless a success. We decided to put the Stanéo digitizer in a box in order to protect it, and we covered it with a piece of plastic. We have then tried to immobilize everything by putting stones here and there.
The weather on the rest of the sol was worrying: there was a lot of wind (impossible thing on Mars because of the low pressure on its surface). I fear that the structure might crumble. Unfortunately, we’ll have to wait until tomorrow to check this out, during the next EVA.

Experiment: Balloon
Person filling in the report: Simon Bouriat

We were first supposed to use the balloon during the EVA this morning. Unfortunately, the weather was not on our side. We still don’t have enough results to make any conclusions. We hope to use it again next week.
The crew engineer and I started working on our new project about emergency procedures. There are three main points in that study. First, we will study the emergency procedures observed during scuba diving. One of the big step in this study is the comparison between this sport’s dangers and EVA’s ones. Secondly, we will try to set up different emergency procedures depending on the issues and brief the rest of the crew about it. Finally, we will simulate unexpected health or material issues during EVAs and try to apply these procedures.

Simon Bouriat – HSO Crew 175

Commander Report – February 17th

Crew 175 Commander Report 17Feb2017

Dear Earth,

Sol 5: back to normal after yesterday’s festivities.

Mouâdh, Victoria, Simon and I went on EVA this morning to install the
seismometer in the hole we dug close to the Main Road, South of the
Hab. We had also planned to deploy the balloon for the second time,
but the wind was too strong. The exploration we did afterward led us
to the top of the Ridge, where we discovered a stunning panorama: in
front of us the red plain enclosed in mud hills, behind us the
colossal mesas of Skyline Ridge. Our hike offered us an unusual sight
on the MDRS from above. We should definitely come back up here on a
later EVA!

A very strong wind rose after lunch, blowing away all the outside
stuff that was not attached (typically: the unused cardboard we stored
yesterday). In order to avoid further mess and to take a look at the
potential damages around the station, Louis Maller and Xavier went out
for a short EVA. In the end, no hurt was noted on the different
facilities, but Louis and Xavier preventively fixed a water tank on
the rear of a rover. We were glad to see that our upgrade on the
tunnel stood still (we call it “The Wall”). This storm reminded us of
the non-realistic beginning of Andy Weir’s novel “The Martian”, with a
minor difference though: the terrestrial wind is strong enough to blow
away heavier things than sand… So better stay aware, even if this
situation is totally impossible under Mars thin atmosphere.

Otherwise, there is nothing to worry about. The crew stays in a good
mood, science goes on and we have plenty of food left (at least for
the moment).

Ad Astra!

Sol Summary – February 17th

Crew 175 Sol Summary 17Feb2017


SOL: 5
Person filling out Report: Louis MALLER, XO
Summary Title: Gone with the wind
Mission Status: Successful EVA, all systems go, work on experiments ongoing
Sol Activity Summary: sport, science work on different experiments, nominal EVA and short unscheduled EVA.
Look Ahead Plan: Tomorrow will see another EVA, and we are looking forwards for the wind to calm down in order to be able to use the balloon again.
Anomalies in work: /
Weather: a bit cloudy, cooler than usual, very windy in the afternoon
Crew Physical Status: Crew feeling well
EVA: EVA #3 successfully deployed the seismometer experiment, and then explored the area around the deployment zone, and realized some tests with talkie range. EVA #4 realized its task and then checked the hab because of the wind, everything was normal except some trash that had flown away.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #5 request
– EVA #3 report
– EVA #4 report
– Science report

EVA Report – February 17th

Crew 175 EVA #4 Report 17Feb2017

EVA Report:

EVA #4

Crew members (2 crewmembers): Xavier Rixhon, Louis Maller

Location: Immediate vicinity of the hab
Vehicles used: /

Departure time:  15:54
Return time: 16:06
Duration: 12min


We exited the airlock under some quite heavy wind. We put the water tank on the rover, which was more difficult than expected as it was heavier than we had thought, difficult to grab and handle, especially wearing spacesuits.
We then described a circle around the hab, got a few trash items that had flown away, checked that everything seemed okay. We put the tunnel extension back in place at the end point which had moved slightly. We didn’t notice any damage anywhere on the MDRS

EVA Report – February 17th

Crew 175 EVA Report 17Feb2017

EVA #3

Crew members: Arthur Lillo (EVA leader), Mouâdh Bouayad, Simon Bouriat
et Victoria Da-Poian (EVA buddies)

•       South of the Hab to bury the seismometer (12N 518500, 4250000)
•       Hab Ridge Road, South of the Hab (12N 518000, 4250100)
•       East of the Hab (12N 518700, 4250700)

Vehicles used: Phobos and Deimos

Departure time:  9:03 AM
Return time: 11:38 AM
Duration: 2h35


•       Around the Hab exploration
•       EVA training for newcomers
•       Seismology experiment: dig a deeper hole and bury the seismometer,
the acquisition system and the battery


The Engineering Check was quick today, no refill was needed. We did
not take the balloon because of the risk of wind. We took the rovers
to go to the plain just North of Camel Ridge, where we dug a hole two
Sols ago. At 9:25 we installed the car battery, the acquisition system
and the seismometer, and covered them with dirt and plastic planks to
protect them from the bad weather. At 9:50 we began our exploration:
we left the seismometer to walk West of its location, where we found
an access to the top of Hab Ridge. The view was absolutely stunning
despite the cloudy sky. We found Hab Ridge Road up there and followed
it until we could see the Hab and establish radio contact with HabCom.
At 10:30, we headed back to the rovers on the same path and lose
contact again. We went back to the Hab’s crossroad and left the rovers
there to explore the Eastern area. We discovered the intriguing site
called Marble Ritual, and enjoyed the view on Pooh’s Corner from the
top of a small hill. After that, we returned to the Hab earlier than